# Semantic Interaction Information mediates compositional generalization in latent space

### 28 Mar 2026 v1

@@ 1 Introduction
To address this shortcoming, we must be able to differentiate meaningful intelligence from that which is trivial. [18] defines intelligence as ’skill-acquisition efficiency’, emphasizing the capacity to build upon prior knowledge. This relies on the ability to recognize or propose an unseen combination of elements as a novel solution, known as compositional generalization. In this paper, we formalize compositional generalization with abstract variables as the elemental units, using the principle of Interaction Information– mutual information among three or more random variables [106]. Interaction information has been used to understand sensory modality integration [73] and causal inference [30], but has not yet been put into a Reinforcement Learning (RL) framework. We bridge this gap by unifying Interaction Information with Semantic Information– the subset of information that is meaningful for survival [50].
To exemplify the relationship between Semantic Information and Interaction Information, consider an agent with two possible actions (i.e. causal interventions), where reward $r$ is determined by an exclusive or (XOR) operation between the agent’s action $a$ and a hidden (i.e. latent) binary variable $c$. XOR demonstrates Interaction Information in its purest form [79]: Action $a$ and reward $r$ are independent unless the agent knows the value of auxiliary variable $c$. We consider such variables as contextual since they mediate the dependencies between action, observations and reward, while not being directly causal. Here, we introduce Semantic Interaction Information (SII)– the contribution of Interaction Information between latent variables (and observables) to inference of the optimal action. More generally, SII emerges when conditioning on observables changes the mutual information between contextual variables and the optimal action. Conceptually, SII measures how much more certain an agent becomes about its goal by accounting for latent variable interactions, rather than treating them in isolation. Moreover, when variables interact in a systematic manner, an agent can use this to generalize to novel combinations of latent variables.
@@ 2.2 Partially Observable Markov Decision Processes
Animals often learn without the need for concrete examples or symbols to reason with. For instance, riding a bike can become automatic without an understanding of all its mechanisms. Such abstract reasoning is often implicit—a ’gut-feeling’ that is effortlessly utilized but not as easily explained. Inference under such uncertainty can be modeled with a POMDP [62, 43]. POMDPs utilize a belief-state, a probability distribution over possible world states which is updated either passively by observation or actively via goal-directed exploration [39, 46]. Distributional representations suffer from the curse of dimensionality, so effective agents must encode goal-relevant information into a compressed approximation of the belief-state [5, 80]. This is consistent with the observation that cortical dynamics often lie on a low dimensional manifold [56, 20]. Furthermore, belief-like representations have even been implicated in the dopaminergic neurons involved in RL [36, 35]. In this work, we consider the compositional generalization that emerges from goal-directed inference in a POMDP. Observations in the Cognitive Gridworld are fundamentally ambiguous, conveying no information without an approximate belief-state which must be learned naturally from reward.
@@ 2.3 Generative modeling
The brain is hypothesized to cluster observations by independent latent causes [29, 25] through mechanisms such as attention and computation through dynamics [65, 103, 102]. Together, these form a traversable cognitive map [100, 8, 28]—an abstract space of task-relevant variables. This capability is believed to emerge via meta-learning [53, 105], wherein agents learn a repertoire of computations that can be reconfigured as needed for flexible inference [111, 101, 11]. Using conditional generative models to produce data for the agent to train on has enabled recent progress in meta-learning. For instance, PEARL learns a posterior over contextual variables that summarize recent experience, and conditions the agent on samples from this posterior to facilitate few-shot acquisition of novel goals [76]. Additionally, [2] show that diffusion-based decision-making can generalize by composing conditioning variables at inference time (albeit in fully observable environments). [1] consider the task of continual learning to be identifying interactions that are shared between datasets. Using this assumption, their VASE algorithm learns abstractions in the data, such as object position, and recombines them to generate augmented data to train on. Such developments illustrate the value of evaluating general intelligence in a simple, standardized setting centered around compositional abstractions. To demonstrate the usefulness of our setting, we show that when interactions occur systematically, the manner in which latent variables interact can be consolidated into embeddings, similar to consolidation by the Hippocampus into long term memory. Learning these compositional embeddings, which encode how a variable will interact with others, requires conditioning a generative model on the inference model, and vice versa. After learning, the chain of representation and classification allows us to invert the process, and evaluate latent states with respect to a preferred observation without need further interactions with the world.
@@ 2.4 Setup
Our goal is to facilitate the study and development of agents that can generalize from goal-directed inference. To simplify our scope, we assume that for any specific goal, only a small number of latent variables are relevant, and that they interact in a systematic and learnable manner. We refer to the variables as states and their values as realizations, which are governed by a random process wherein each latent state, in a context of $C$ relevant states, takes on a value from $R$ possibilities:
$\mathbf{r}\overset{\text{i.i.d.}}{\sim}U(\{1,\dots,R\})^{C},\quad\text{where }r_{c}\in\mathbf{r}\text{ is the realization of state }c.$
The agent observes the world through a $d_{o}$-dimensional observation space $\mathcal{O}$ (Figure 1), where the probability of each independent observable $o^{i}\in\{0,1\},\forall o^{i}\in\mathbf{o}$ depends on $\mathbf{r}$ via a joint likelihood parameterized by state interactions, $Z$. For simplicity and learnability, we assume that interactions are pairwise and independent, resulting in $|Z|=d_{o}\times C\times(C-1)$ interactions. The likelihood of observing $o^{i}=1$, denoted by the shorthand $\ell_{\mathbf{z}^{i}}(\mathbf{r}),\quad\forall\mathbf{z}^{i}\in Z$ (where $\ell_{\mathbf{z}^{i}}(\mathbf{r})\equiv P_{\mathbf{z}^{i}}(o^{i}=1\mid\mathbf{r})$) is given by a tensor of order $C$ with $R$ entries along each axis. Each interaction $z^{i}_{cc^{\prime}}\in\mathbf{z}^{i}$ (denoting the modulation from state $c^{\prime}$ onto state $c$) shifts the phase of the periodic (w.r.t to $r_{c}$) likelihood. This functionally ”rolls” values of $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ along axis $c$. Since our goal is to generate Interaction Information in an interpretable manner, $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ is modeled around the exclusive or operation, while remaining adjustable to other stationary POMDPs. Further details can be found in subsection A.2.
Figure 1: Environment schematic for $C=2$. Observations are generated stochastically. State interactions parameterize the likelihood and state realizations fix the probability of each observable.
An episode $E$ is defined by the tuple $(\mathbf{r},g,Z,\mathbf{o}_{1:T})$, containing the realizations, a randomly selected goal index $g\sim\{1,\dots,C\}$, the set of interactions, and a trajectory of $T$ i.i.d. observations. The agent’s objective is to infer the realization of the goal state, $a^{\star}=r_{g}$, and select the corresponding action $a\in\{1,\dots,R\}$. The agent can estimate $r_{g}$ by using sequential Bayesian inference:
@@
where the joint likelihood of the independent observables follows a geometric distribution:
$P_{Z}(\mathbf{o}=\mathbf{o}_{t}\mid\mathbf{r})=\prod_{i=1}^{d_{o}}\ell_{\mathbf{z}^{i}}(\mathbf{r})^{o_{t}^{i}}(1-\ell_{\mathbf{z}^{i}}(\mathbf{r}))^{1-o_{t}^{i}}.$
@@ 3 Results
A natural baseline for our setting is a Bayesian observer whose marginal beliefs ($B_{tc}$) are updated independently. The likelihood for this Naive observer is acquired by marginalizing the joint likelihood (Figure 2a). Semantic Interaction Information quantifies the contribution of Interaction Information to marginal beliefs as the kullback-leibler divergence ($\mathcal{D}_{\mathrm{KL}}$) between $B_{tg}$ acquired through Joint and independent inference:
@@
(A formal derivation is provided in subsection A.3). Figure 2b compares the posterior distributions of the Bayesian observers. We observe that, for both Joint and Naive Bayes, the average performance (calculated as $\langle 1-[\arg\max_{r}\bigl(B_{tgr}\bigr)=a^{\star}]\rangle_{E}$ using Iverson bracket notation) improves with the number of evidences received and declines with the number of interactions. However, the relative accuracy attained by joint inference when compared to independent inference grows rapidly with the number of interactions, in accordance with the Semantic Interaction Information. In summary, the Cognitive Gridworld demonstrates that even in stationary POMDPs, systematic gaps emerge between nuanced and Naive Bayesian observers, despite both possessing perfect memory.
Figure 2: The cost of Naive Bayes grows with time and interactions. (a) Joint (matrices) and independent (vectors) likelihoods for example observables. (b) Top: Accuracy of Joint (left) and Naive (right) Bayes across varying context sizes. Bottom: Relative accuracy (left) and Semantic Interaction Information (right) of Joint versus Naive Bayes. Circles mark four equidistant reference time-points throughout inference.
@@ 3.2 Experiment 1: The Cognitive Gridworld requires flexible neural dynamics
The Cognitive Gridworld inverts the standard dependency structure of most tasks. Typically, independent features of an observation space (e.g., pixels or tokens) become conditionally dependent given the realization of a latent variable (e.g., a label or concept). In a CG, the latent variables become conditionally dependent given the observations. The lack of spatial or temporal correlations in the observation space yields a uniform distribution of data, such that all learnable structure is latent. The network performing sequential inference, referred to as the Classifier, consists of Long-Short-Term-Memory (LSTM) units and receives $Z$ and $\mathbf{o}_{1:T}$. To demonstrate that arbitrary neural dynamics that encode $Z$ and $\mathbf{o}_{1:T}$ are insufficient, we benchmark a Fully Trained RNN against reservoir computing (Echo State RNN). While Echo State dynamics provide random temporal features on which to learn a function of recent inputs, they suffer from fundamental limitations that may be viewed as a tradeoff between discrimination and generalization, memory and nonlinearity or recall and precision [3, 21, 6, 40, 98, 97]. As a result, we predicted that the performance of an Echo State network would resemble that of a Naive Bayesian observer.
@@ 3.2.1 Interaction Information underlies network capability at inference time
To ensure the Echo State network has a rich teaching signal to learn from, we train with the optimal belief as the target (see subsection 5.1 for details), and to ensure performance does not reflect memory limitations, the network outputs belief updates $\Delta B_{t}\in\mathbb{R}^{C\times R}$ which we accumulate over time (Figure 3a). In Figure 3b we observe that the performance of the Fully Trained network is explained by Joint Bayes, meanwhile, as predicted, the Echo State network reaches exactly the accuracy of Naive Bayes (Figure 3c). Crucially, the Echo State network performs optimally in the single latent-state case ($C=1$), where joint inference is equivalent to independent inference. This indicates the sub-optimality at $C=2$ is not due to basic memory limitations, but a specific inability to approximate joint inference. Moreover, the relative accuracy between the Fully Trained and Echo State networks mirrors the exact ratio seen between Joint and Naive Bayes (Figure 3d). These results suggest that task-aligned dynamics [89] are particularly necessary when latent variables interact.
Figure 3: Recurrent Neural Networks align with theoretical predictions. (a) Architecture (left) and gradient flow (right) of the Classifier. Only the goal belief-state receives a gradient. (b-d) Same as Figure 2b (for $C=1,2$) with Fully Trained and Echo State Networks.
@@ 3.3.1 Naturalistic feedback is sufficient for learning interactions
Inspired by the V-JEPA 2-AC architecture [4], we separate the problem into embedding and prediction. Our Representation Classification Chains (RCC) (Figure 6b) use a variational approach where modules bootstrap off each other without sharing gradients: a Classifier predicts state realizations from $\hat{Z}$, while a Generator maximizes the likelihood of observations conditioned on $\hat{Z}$ and $\hat{\mathbf{r}}$ (see subsection 5.2 for details). To train the Classifier we treat the goal belief as a policy. At the end of a trajectory we sample an action and use a cross-entropy-style loss to provide naturalistic feedback that is sparse (end-of-trajectory), partial (goal-state only), and binary (reward/punishment). Concurrently, the Generator network is trained via a self-supervised loss and the gradient flows from $\hat{Z}$ to $\hat{\mathcal{E}}$, teaching the embedding space to be useful for the Generator. Importantly, we define a subset of testing states which, during training, (i) never appear as goals and (ii) never co-occur in a context. Additionally, training contexts contain (at most) one testing state, while testing contexts contain only testing states. Because contextual belief is never evaluated during training, testing states’ embeddings must be learned via their implicit relationships to training goals. Figure 6c illustrates the improvement in test accuracy, throughout training, at each step of inference, demonstrating that $\hat{\mathcal{E}}$ is indeed learnable from naturalistic feedback.
Figure 5: Goals enable generalization from experience through compositional embeddings. Schematic demonstrating compositional generalization to a novel combination of states.
Figure 6: A variational architecture learns compositional embeddings from reward. (a) Active states interact via learnable embeddings to form interactions. (b) Forward pass and gradient flow of the Classifier and Generator. The Classifier learns from rewards while the Generator uses self-supervised-learning (SSL). (c) Testing episode accuracy of the Classifier throughout training.

### 20 Apr 2026 v3

@@ 1 Introduction
To address this shortcoming, we must be able to differentiate meaningful intelligence from that which is trivial. [18] defines intelligence as ’skill-acquisition efficiency’, emphasizing the capacity to build upon prior knowledge. This relies on the ability to recognize or propose an unseen combination of elements as a novel solution, known as compositional generalization. In this paper, we formalize compositional generalization with abstract variables as the elemental units, using the principle of Interaction Information– mutual information among three or more random variables [108]. Interaction information has been used to understand sensory modality integration [74] and causal inference [30], but has not yet been put into a Reinforcement Learning (RL) framework. We bridge this gap by unifying Interaction Information with Semantic Information– the subset of information that is meaningful for survival [50]. To exemplify the relationship between Semantic Information and Interaction Information, consider an agent with two possible actions (i.e. causal interventions), where reward $r$ is determined by an exclusive or (XOR) operation between the agent’s action $a$ and a hidden (i.e. latent) binary variable $c$. XOR demonstrates Interaction Information in its purest form [80]: Action $a$ and reward $r$ are independent unless the agent knows the value of auxiliary variable $c$. We consider such variables as contextual since they mediate the dependencies between action, observations and reward, while not being directly causal. Here, we introduce Semantic Interaction Information (SII)– the contribution of Interaction Information between latent variables (and observables) to inference of the optimal action. More generally, SII emerges when conditioning on observables changes the mutual information between contextual variables and the optimal action. Conceptually, SII measures how much more certain an agent becomes about its goal by accounting for latent variable interactions, rather than treating them in isolation. Moreover, when variables interact in a systematic manner, learning can extrapolate to novel combinations of latent variables.
@@ 2.2 Partially Observable Markov Decision Processes
Animals often learn without the need for concrete examples or symbols to reason with. For instance, riding a bike can become automatic without an understanding of all its mechanisms. Such abstract reasoning is often implicit—a ’gut-feeling’ that is effortlessly utilized but not as easily explained. Inference under such uncertainty can be modeled with a POMDP [63, 43]. POMDPs utilize a belief-state, a probability distribution over possible world states which is updated either passively by observation or actively via goal-directed exploration [39, 46]. Distributional representations suffer from the curse of dimensionality, so effective agents must encode goal-relevant information into a compressed approximation of the belief-state [5, 81]. Incidentally, belief-like representations have been implicated in the dopaminergic neurons involved in RL [36, 35]. Furthermore, cortical dynamics are often found to lie on a low dimensional manifold [56, 20]. In this work, we consider the compositional generalization that emerges from the dynamics of approximate inference. Observations in the Cognitive Gridworld are fundamentally ambiguous, without any inherent spatiotemporal structure that could circumvent the need to learn to approximate the underlying world model.
@@ 2.3 Generative modeling
The brain is hypothesized to cluster observations by independent latent causes [29, 25] through mechanisms such as attention and computation through dynamics [66, 105, 103]. Together, these form a traversable cognitive map [101, 8, 28]—an abstract space of task-relevant variables. This capability is believed to emerge via meta-learning [53, 107], wherein agents learn a repertoire of computations that can be reconfigured as needed for flexible inference [113, 102, 11]. Modern meta-learning algorithms utilize conditional generative models to produce data for the agent to train on. For instance, PEARL learns a posterior over contextual variables that summarize recent experience, and conditions the agent on samples from this posterior to facilitate few-shot acquisition of novel goals [77]. Additionally, [2] show that diffusion-based decision-making can generalize by combining conditioning variables at inference time. [1] consider the task of continual learning to be identifying generative factors that are shared between datasets. Using this assumption, their VASE algorithm learns independent factors in the data and recombines them to generate augmented data to train on. Such research illustrates the value of a simple, standardized meta-learning task centered around compositional abstractions. To demonstrate the usefulness of the Cognitive Gridworld, we show that the manner by which latent variables interact can be consolidated into embeddings. Learning these compositional embeddings, which encode how a variable will interact with others, requires conditioning a generative model on the inference model, and vice versa. To achieve this, we develop Representation Classification Chains and show that, after learning, the process can be inverted to evaluate locations on a novel cognitive map, with respect to preferred observations.
@@ 2.4 Setup
Our goal is to facilitate the study and development of agents that can generalize from goal-directed inference. To simplify our scope, we assume that for any specific goal, only a small number of latent variables are relevant, and that they interact in a systematic and learnable manner. We refer to the actual value of a latent variable as its realization, which is governed by a random process wherein each variable, in a context of $C$ relevant variables, takes on a value from $R$ possible realizations:
$\mathbf{r}\overset{\text{i.i.d.}}{\sim}U(\{1,\dots,R\})^{C},\quad\text{where }r_{c}\in\mathbf{r}\text{ is the realization of variable }c.$
The agent observes the world through a $d_{o}$-dimensional observation space $\mathcal{O}$ (Figure 1). At time $t$, the probability of each independent observable $o^{i}\in\{0,1\}$ within observation $\mathbf{o}_{t}$ depends on realizations $\mathbf{r}$ and is parameterized by variable interactions $Z$. In total, there are $|Z|=d_{o}\times C\times(C-1)$ pairwise and independent interactions. We let $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ denote the likelihood $P_{\mathbf{z}^{i}}(o^{i}=1\mid\mathbf{r})$, which forms a tensor of order $C$ with $R$ entries along each axis. Each interaction $z^{i}_{cc^{\prime}}\in\mathbf{z}^{i}$ (the modulation from variable $c^{\prime}$ onto $c$) shifts the phase of the likelihood with respect to $r_{c}$, functionally ”rolling” values of $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ along axis $c$ (see supplementary Figure 9). To generate Interaction Information in an interpretable manner, we model $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ around the exclusive or operation via a procedure detailed in subsection A.1 (customize-able to any arbitrary function of $Z$). Importantly, these joint likelihoods cannot be factorized without losing information. Finally, an episode $E$ is defined by the tuple $(\mathbf{r},g,Z,\mathbf{o}_{1:T})$, comprising the realizations, a random goal index $g\sim\{1,\dots,C\}$, the interactions, and a trajectory of $T$ i.i.d. observations.
The agent’s objective is to infer the realization of goal variable $r_{g}$ and select optimal action $a^{\star}=r_{g}$ from $a\in\{1,\dots,R\}$. Optimal inference of $r_{g}$ can be achieved with sequential Bayesian updating:
@@
where the joint likelihood of all (independent) observables follows a geometric distribution:
$P_{Z}(\mathbf{o}_{t}\mid\mathbf{r})=\prod_{i=1}^{d_{o}}\ell_{\mathbf{z}^{i}}(\mathbf{r})^{o_{t}^{i}}(1-\ell_{\mathbf{z}^{i}}(\mathbf{r}))^{1-o_{t}^{i}}.$
Figure 1: Environment schematic for $C=2$. Observations $\mathbf{o}_{t}$ are generated stochastically. variable interactions $Z$ parameterize the likelihood $P_{Z}(\mathbf{o}\mid\mathbf{r})$ and variable realizations $(r_{1},r_{2})$ fix the probability of sampling each observable $o^{i}$.
Thus, while the real world involves a vast number of latent variables, the agent’s goal effectively reduces the world to a context of $C$ relevant variables—specifically, the goal variable itself and the contextual variables that generate SII. In other words, although the state space may be high dimensional, the optimal policy likely lies on a smaller $C$-dimensional manifold at any given time.
@@ 3 Results
A natural baseline for our setting is a Bayesian observer whose marginal beliefs ($B_{tc}$) are updated independently. The likelihood for this Naive observer is acquired by marginalizing the joint likelihood (Figure 2a). Semantic Interaction Information quantifies the contribution of Interaction Information to marginal beliefs as the kullback-leibler divergence ($\mathcal{D}_{\mathrm{KL}}$) between $B_{tg}$ acquired through Joint and Naive Bayesian inference:
@@
(A formal derivation is provided in subsection A.2). Figure 2b compares the posterior distributions of the Bayesian observers. We observe that, for both Joint and Naive Bayes, the average performance (calculated as $\langle 1-[\arg\max_{r}\bigl(B_{tgr}\bigr)=a^{\star}]\rangle_{E}$ using Iverson bracket notation) improves with the number of evidences received and declines with the number of interactions. However, the relative accuracy attained by joint inference when compared to naive inference grows rapidly with the number of interactions, in accordance with the Semantic Interaction Information. In summary, the Cognitive Gridworld demonstrates that even in stationary POMDPs, systematic gaps emerge between nuanced and Naive Bayesian observers, despite both possessing perfect memory.
Figure 2: The cost of Naive Bayes grows with time and interactions. (a) Example Joint (matrices) and marginalized (vectors) likelihoods. (b) Top: Accuracy of Joint (left) and Naive (right) Bayes across varying context sizes. Bottom: Relative accuracy (left) and Semantic Interaction Information (right). Circles mark four equidistant reference time-points throughout inference.
@@ 3.2 Experiment 1: The Cognitive Gridworld requires flexible neural dynamics
The Cognitive Gridworld inverts the standard dependency structure of most tasks. Typically, independent features of an observation space (e.g., pixels or tokens) become conditionally dependent given the realization of a latent variable (e.g., a label or concept). In a CG, conditioning on latent variable realizations causes observables to become independent. The lack of spatial or temporal correlations in the observation space yields a uniform distribution of data, such that all learnable structure is latent. The network performing sequential inference, referred to as the Classifier, consists of Long-Short-Term-Memory (LSTM) units and receives $Z$ and $\mathbf{o}_{1:T}$. To demonstrate that arbitrary neural dynamics that encode $Z$ and $\mathbf{o}_{1:T}$ are insufficient, we benchmark a Fully Trained RNN against reservoir computing (Echo State RNN). While Echo State dynamics provide random temporal features on which to learn a function of recent inputs, they suffer from fundamental limitations that may be viewed as a tradeoff between discrimination and generalization, memory and nonlinearity or recall and precision [3, 21, 6, 40, 99, 98]. As a result, we predicted that the performance of an Echo State network would resemble that of a Naive Bayesian observer.
@@ 3.2.1 Interaction Information underlies network capability at inference time
To ensure the Echo State network has a rich teaching signal to learn from, we train with the optimal belief as the target (see subsection 5.1 for details), and to ensure performance does not reflect memory limitations, the network outputs belief updates $\Delta B_{t}\in\mathbb{R}^{C\times R}$ which we accumulate over time (Figure 3a). In Figure 3b we observe that the performance of the Fully Trained network is explained by Joint Bayes, meanwhile, as predicted, the Echo State network reaches exactly the accuracy of Naive Bayes (Figure 3c). Crucially, the Echo State network performs optimally in the case of a single latent variable ($C=1$), where joint inference is equivalent to independent inference. This indicates the sub-optimality at $C=2$ is not due to basic memory limitations, but a specific inability to approximate joint inference. Moreover, the relative accuracy between the Fully Trained and Echo State networks mirrors the exact ratio seen between Joint and Naive Bayes (Figure 3d). These results suggest that task-aligned dynamics [90] are particularly necessary when latent variables interact. A comparison with $C=3$ can be seen in supplementary Figure 16a.
Figure 3: Recurrent Neural Networks align with theoretical predictions. (a) Architecture (left) and gradient flow (right) of the Classifier. Only the goal belief-state receives a gradient. (b-d) Same as Figure 2b (for $C=1,2$) with Fully Trained and Echo State Networks. Markers indicate four equidistant reference time-points throughout inference.
@@ 3.3 Experiment 2: Learning interactions requires variational inference
Thus far, we have established how Interaction Information can impede optimal action selection. While networks in Experiment 1 were trained with explicit access to the true interactions $Z$, subsequent experiments assume these latent interactions are unknown but learnable. To achieve this, we introduce compositional embeddings that encode how variables interact. Consider a large set $\mathcal{S}$ of latent variables, where each individual variable (indexed by $s\in\mathcal{S}$) represents a subset of mutually exclusive realizations of the world. These may be the seasons or the time of day, for example. For a specific goal, like waking up on time, and given the observed sunlight, most other latent variables are irrelevant. However, if the interaction between season and time correlates with how the season would interact with another latent variable, possibly relevant to a future goal, this relationship is encoded in season’s compositional embeddings. Formally, we associate variable $s$ with a set of learnable embeddings $\{(\mathbf{k}_{s}^{i},\mathbf{q}_{s}^{i})\}_{i=1}^{d_{o}}\in\mathcal{E}$. The context, i.e. subset of relevant variables $C\subset\mathcal{S}$, has interactions $z_{cc^{\prime}}^{i}=(\mathbf{q}_{c}^{i})^{\top}\mathbf{k}_{c^{\prime}}^{i}$ for all distinct pairs $c,c^{\prime}\in C$ (Figure 5).
Figure 5: Compositional embeddings are learned indirectly via goals. Schematic demonstration of compositional generalization in latent space. Training episodes (top) contain at most one testing variable, which is never the goal (green). Testing episodes (bottom) consist entirely of testing variables. Success requires testing variable embeddings to be learned through their implicit relationships to training goals. This setup evaluates zero-shot generalization to novel contexts and novel goals.
@@ 3.3.1 Naturalistic feedback is sufficient for learning interactions
Inspired by the V-JEPA 2-AC architecture [4], we separate the problem into embedding and prediction. Modules of our Representation Classification Chains (RCCs) (Figure 6b) bootstrap off each other without sharing gradients: a Classifier uses $\hat{Z}$ to predict realizations $\hat{\mathbf{r}}$, while a Generator uses $\hat{\mathbf{r}}$ to predict observations (see subsection 5.2). To train the Classifier we treat the goal belief as a policy, sampling a single action and using a cross-entropy-style loss to provide feedback that is sparse (sampled at the end of the trajectory), partial (goal-variable only), and binary (reward/punishment). Concurrently, the Generator network is trained via a self-supervised loss and the gradient flows from $\hat{Z}$ to $\hat{\mathcal{E}}$, teaching the embedding space to be useful for prediction. Figure 6c illustrates testing episode accuracy throughout training, demonstrating that $\hat{\mathcal{E}}$ is indeed learnable from naturalistic feedback.
Figure 6: A variational architecture learns compositional embeddings from reward. (a) Relevant variables interact via learnable embeddings to form interactions. (b) Forward pass and gradient flow of the Classifier and Generator. The Classifier learns from rewards while the Generator uses self-supervised-learning (SSL). (c) Testing episode accuracy of the Classifier throughout training.